Question 1

During 2018, the following events occurred for Clouseau Consultants

  1. Clouseau provided $35,744 of consulting services to its clients but no cash has been collected.
  2. Clouseau collected $21,274 cash from customers in partial settlement of its accounts receivable.
  3. Clouseau paid the instructor $11,567 cash for teaching training courses (salary expense).
  4. Clouseau paid $3,650 for supplies. Half of the supplies were used in 2018.
  5. Clouseau paid $4,484 in dividends.

After recording all these transactions, Clouseau’s total change in stockholders' equity will be $_________

Question 2

Bailey Co earns $16,205 of revenue on account and in $5,521 cash revenue transactions in Year 1. Cash collections of receivables amount to $7,624 in Year 1 with the remainder being collected in Year 2. Based on this information alone the company’s financial statements would show

Total Revenue in Year 1 of $________________
Answer #1

Part 1

After recording all these transactions, Clouseau’s total change in stockholders' equity will be $17868

Clouseau’s total change in stockholders' equity = increase in revenue – increase in expenses – dividends paid

Clouseau’s total change in stockholders' equity = 35744 – (11567+(3650/2))-4484 = $17868

Clouseau’s total stockholders' equity is increases by $17868.

Part 2

Total Revenue in Year 1 of $21726

Revenue in Year 1 = credit revenue + cash revenue = 16205 + 5521 = $21726